OpenStudy (xochital):

What is the slope of a line that passes through the point (−5, 3) and is parallel to a line that passes through (2, 13) and (−4, −11)?

OpenStudy (anonymous):

just wait for a while

OpenStudy (anonymous):

Parallel means the two slopes are the same. So, find the slope of the 2 given points, (2 , 13) and (-4 , -11) using the formula. Slope = (y2 - y1) / (x2 - x1) where (x1 , y1) and (x2 , y2) are the 2 points.

OpenStudy (anonymous):

Since parallel lines have equal slopes Slope of the required line = slope of the line passing through (2, 13) and (−4, −11) =\[=\frac{y_1-y_2}{x_1-x_2}=\frac{13-(-11)}{2-(-4)}=\frac{13+11}{2+4}=\frac{24}{6}=4\]

OpenStudy (anonymous):

Hence, the slope of a line that passes through the point (−5, 3) is 4

OpenStudy (xochital):

thank you so much :)